Working together, Adam and Chris can reseal a driveway in
6 hours. Working alone, Adam can reseal the driveway in 15
hours. How long would it take Chris, working alone,to reseal the
driveway?

Answer by mananth(16946)

Working together, Adam and Chris can reseal a driveway in
6 hours.
Together they can reseal 1/6 of the driveway in 1 hour
Working alone, Adam can reseal the driveway in 15
Adam can reseal1/15 of the driveway in 1 hour
Let Chris take x hours to complete the job alone
He does 1/x of the job in 1 hour
1/15 + 1/x = 1/6
1/x = 1/6 -1/15
1/x = 9/90 = 1/10
x= 10
Chris take 10 hours to complete the job alone